can anyone show me a table and then solve for the following word problem? If you drive for 2 hours at 80 km/hr, how fast must you rive during the next hour in order to have an average speed of 75 km/hr
Answer by mananth(16946)

can anyone show me a table and then solve for the following word problem? If you drive for 2 hours at 80 km/hr, how fast must you rive during the next hour in order to have an average speed of 75 km/hr
...
hours speed distance
2-------80---------160
..
1--------x----------1*x
...
Total average = total distance/total time
total distance = 160+x
total time = 3 hours
average required = 75 mph
...
(160+x)/3 = 75
(160+x)=3*75
160+x=225
x=225-160
x=65 mph
he should drive at 65 mph
...
m.ananth@hotmail.ca